OpenStudy (anonymous):

Miss Crystal Palace just purchased – you guessed it – a crystal palace for $650,000. She makes a 25% down payment and finances the balance with a mortgage at an interest rate of 5.6% for 15 years, making monthly payments. a. What amount will she borrow? b. Complete the amortization table for the first three payments. (Show all work!)

OpenStudy (phi):

can you answer (a) ? 75% of 650,000 or 650,000- 25% of 650,000

OpenStudy (anonymous):

162500?

OpenStudy (anonymous):

oh wait no. 649999.75?

OpenStudy (anonymous):

no isn't my first answer right?

OpenStudy (phi):

If you put down a down-payment, do you expect it to be all 650,000? to do this, remember that 25% is the same as 25/100 (or 1/4)

OpenStudy (anonymous):

is it 162500?

OpenStudy (phi):

that is the down payment. How much is the loan?

OpenStudy (anonymous):

162500(1+0.056/12)^12*15 162500(1+0.056/12)^180 (162500)(2.311845392)=$375,674.88

OpenStudy (anonymous):

that's what I did...

OpenStudy (anonymous):

ohh in my equation where i put 162500.. is that supposed to be 487500 instead?

OpenStudy (phi):

162,500 is 1/4 of the 650,000 162,500 is the amount paid up front. You must now borrow the remainder. so your work up above looks good except 162500 is not the correct amount for the loan

OpenStudy (anonymous):

the amount of the loan is 1,127,024.63?

OpenStudy (phi):

For (a), I think they want 487,500 to do (b) you have to figure out how much the total loan ends up costing. Is that the 1,127,024 number?

OpenStudy (anonymous):

yes the total loan I calculated is 1,127,024.63 I believe.

OpenStudy (phi):

now find the monthly payment

OpenStudy (anonymous):

how? what formula?

OpenStudy (anonymous):

6261.25 a month?

OpenStudy (phi):

now you can do the amortization schedule start with the loan balance at the start 487500 add one month of interest (0.056/12)*487500 then subtract off the payment to get the loan balance for the next month. do the same thing for the 2nd (and 3rd) month, except use the new loan principal

OpenStudy (anonymous):

what new loan principal?

OpenStudy (phi):

when you make a payment, the amount you owe goes down

OpenStudy (anonymous):

(0.056/12)*487500= 2275

OpenStudy (anonymous):

487500-2275= 485225 for the next month?

OpenStudy (anonymous):

2264.38--485235.62 2264.43--482971.19

OpenStudy (anonymous):

how do I put that in a chart though with the.. amount of payment interest payment applied to principal and balance?

OpenStudy (phi):

OK, I just checked about loans. there is a formula to figure out the monthly payment FV= payment *( (1+i)^n -1)/i where future value FV is the number you found, does this look familiar?

OpenStudy (anonymous):

yeah, give me a sec.

OpenStudy (anonymous):

so the formula I used gave me 9268.65 as the monthly payment

OpenStudy (phi):

how? I get something else

OpenStudy (anonymous):

P=1127024.63 n=180 r/m=0.56/12=0.0046666667 1127024.63(1+0.0046666667)^180=R[(1+0.0046666667)^180-1)/(0.0046666667)] 2605506.713=R(281.1097278) R=9268.65

OpenStudy (anonymous):

the amount is 1127024.63 the annual rate is 5.6% and the time is 15 years

OpenStudy (anonymous):

and the interest payment is 5259.45

OpenStudy (phi):

I match up, but with 1127024.63=R(281.1097278)

OpenStudy (anonymous):

4009.2 is the applied principal and the balance is 1127024.63

OpenStudy (anonymous):

do I not solve 1127024.63(1+0.0046666667)^180 that?

OpenStudy (anonymous):

because in this example I have is shows to do that..

OpenStudy (phi):

for your formula, it looks like you should use the original loan amount on the left hand side. Do you notice that the left side is future value (1,237,024) which can be calculated as 487500*(1+0.056/12)^180

OpenStudy (anonymous):

ohh so after solving it becomes 1127024.63

OpenStudy (anonymous):

so is the monthly payment 4009.20?

OpenStudy (phi):

yes. so now to the amort. to answer amount of payment interest payment applied to principal and balance? you have the first part. interest is principal* 0.0046666667 (5.6% /12) so you start with P (487500) add interest subtract payment get new P (call it P2 for second month) The interest payment is the amount of interest due. the applied to principal is the amount of the payment left over after paying the interest P2 is the balance

OpenStudy (anonymous):

wait wait.. is the interest payment 5259.45? (1127024.63)(0.0046666667)(1)

OpenStudy (phi):

No, you use Present value (not future value) when you apply the interest.

OpenStudy (anonymous):

(487500)(0.0046666667)(1)=2275.00

OpenStudy (phi):

that looks good. now the payment pays that interest and what is left over pays down the principal

OpenStudy (anonymous):

4009.20-2275.00=1734.20?

OpenStudy (phi):

so now how much is the new principal?

OpenStudy (anonymous):

485765.8?

OpenStudy (phi):

so put those numbers in the first row. now do it again with P=485765.80

OpenStudy (anonymous):

so.. 1734.20*0.0046666667*1?

OpenStudy (phi):

You seem to be missing the main point. you pay interest on the dollars you owe (think of it as rent) after one payment you owe 485765.80. that is what the interest is on (not the 1734)

OpenStudy (anonymous):

2266.91? that is the next interest payment for month 2?

OpenStudy (phi):

yes

OpenStudy (anonymous):

then 1742.29 is the applied principal for month 2?

OpenStudy (phi):

yes

OpenStudy (anonymous):

then the next balance is 484023.51?

OpenStudy (phi):

yes. one more payment to go

OpenStudy (anonymous):

2258.78 for the interest payment?

OpenStudy (anonymous):

and 1750.42 and the applied principal?

OpenStudy (phi):

yes

OpenStudy (phi):

Here is a site that computes the amort http://bretwhissel.net/cgi-bin/amortize

OpenStudy (anonymous):

thank you so much for your help!

OpenStudy (phi):

It helped remind me how to do this stuff. It is a bit complicated.

OpenStudy (anonymous):

haha yeah, well before I had absolute no idea where to start. Thank you so much for helping me..